OpenStudy (anonymous):

What is 40/32 in simplest form?

OpenStudy (anonymous):

8

OpenStudy (shamil98):

\[\large \frac{ 40 }{ 32 } ==> \frac{ 20 }{ 16 } ==> \frac{ 10 }{ 8 } ==>\frac{ 5 }{ 4 }\]

OpenStudy (unklerhaukus):

shamil98 has canceled factors of 2